# Decentralization of Ethereum’s Builder Market

### v1

@@ C.1. Importance of private order flows
To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions. We first identify all private order flows using methods described in Sec. 3.1. Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees (Ethereum, 2024) and direct transfers (Flashbots, 2024a) to the block builder) to the total values extracted from all transactions in the block.
C.2. Case Study for Maestro
@@
To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig. 12. Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig. 4, i.e., from May to November, Maestro’s contribution accounts for approximately 80%! This means Maestro used to be an important upstream of MEV-Share and MEV Blocker. In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
@@
Figure 12. Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
C.3. Requirements of Providers
@@
Telegram trading bot Banana Gun (Banana Gun, 2023), Maestro (Maestro, 2023a), Unibot (Unibot X, 2023) The requirements are the same as the channels they use or unclear.
Channel BloXroute (bloXroute Labs, 2024) Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
Merkle (Merkle Software Inc., 2023) Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
MeowRPC (Meow RPC, 2024) The transactions are sent to the top builders which MeowRPC’s team trusts.
MEV Blocker (MEV Blocker, 2023) Builders must maintain at least 1% market share over one week.
MEV-Share (Flashbots, 2024) Builders must demonstrate a competitive market share for consideration.
Transaction Boost (Blocknative, 2024) Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
@@
Table 6. Collections of the private order flow providers and their requirements.
C.4. Minimal Cost for 1% Market Share
We now present how we compute the minimal subsidy for a builder to account for 1% market share over one week. We assume that all builders have access to public order flows but the new builder cannot access any of the private order flows. Furthermore, this new builder can win the auction if their bid value is the same as other builders. Other builders will not submit bids beyond their true value, defined as the total revenue of their blocks. Or they will not increase subsidies if they already subsidize.
@@ C.4. Minimal Cost for 1% Market Share
Appendix D Competition and Efficiency in the MEV-Boost auctions
D.1. Measuring block-building capabilities
Figure 13. Violin graphs for the distribution of the CV across three MEV tiers, for top, middle, and tail builders. The CV measures the inequality of block-build capacity across builders.
D.2. Implications of uncompetitive auctions
Figure 14. The distribution of the CIs of MEV-Boost auctions

### v3

@@ F.1 Importance of Private Order Flows
To quantify the importance of private order flows for building valuable blocks, we compute the ratio of block values derived from private transactions. We first identify all private order flows using methods described in Sec. 3.1. Then, for each block, we compute the ratio between values extracted from private transactions (i.e., the total transaction priority fees [21] and direct transfers [32] to the block builder) to the total values extracted from all transactions in the block.
F.2 Pivotal Private Order Flow Providers
Third, we observe a notable downtrend in the pivotal level of MEV-Share and MEV Blocker starting from November 2023, accompanied by an uptrend in Maestro, a popular Telegram trading bot [55], beginning in December 2023. This is because Maestro used to be an important upstream of MEV-Share and MEV Blocker but stopped in November 2023, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
Although platforms like MEV-Share and MEV Blocker positively affect decentralization, the incentives of their upstream providers may not be aligned. For example, Mastro may prefer to send flows to builders directly for a better price than what MEV-Share and MEV Blocker can offer. It is surprising to see that upstream providers can significantly swing the significance of platforms.
Finding V: We identified five pivotal providers (MEV-Share, MEV Blocker, jaredfromsubway.eth, Banana Gun, Maestro) who had a sustained influence on the outcome of more than half of MEV-Boost auctions. A positive trend towards decentralization is observed as the pivotal level of the searcher jaredfromsubway.eth decreased, while the pivotal level of MEV-Share and MEV Blocker increased. However, their reliance on upstream providers like Maestro is alarming.
@@ F.2 Pivotal Private Order Flow Providers
Case study for maestro.    To delve deeper into the relationship between these two trends, we plot Maestro’s contribution to the total daily value of transactions sent through MEV-Share and MEV Blocker during our study in Fig. 14. Surprisingly, during the time periods when MEV-Share and MEV Blocker were highly pivotal in Fig. 8, i.e., from May to November, Maestro’s contribution accounts for approximately 80%! This means Maestro used to be an important upstream of MEV-Share and MEV Blocker. In November 2023, Maestro stopped sending transactions to MEV-Share and MEV Blocker, causing the pivotal levels of MEV-Share and MEV Blocker to drop significantly.
@@
Figure 14: Daily percentage of the value of the transactions originating from Maestro for both MEV-Share and MEV Blocker.
Figure 15: Weekly builder-specific pivotal levels for the top-3 builders for slots with winning bid values of no less than 0.5 ETH.
F.3 New Integration Pattern
To illustrate the new integration pattern between providers and builders, we consider the MEV significant in a slot if its winning bid value is no less than 0.5 ETH. Additionally, we focus only on the provider that is the sole pivotal provider in a slot, meaning this provider provides significant MEV to the builder, enabling the builder to win the auction. We apply the above adjustments to builder-specific pivotal levels, and the results are shown in Fig. 15. Compared to Fig. 9, we observe that Banana Gun and Titan, as well as jaredfromsubway.eth and Beaver, follow this integration pattern: a provider selectively shares valuable order flows with specific builders when MEV is significant.
F.4 Requirements of Providers
@@
Telegram trading bot Banana Gun [5], Maestro [55], Unibot [82] The requirements are the same as the channels they use or unclear.
Channel BloXroute [9] Builders are granted access to different tiers, each providing varying privileges and permissions, based on an assessment.
Merkle [59] Builders must have a proven track record of accurately and honestly computing bundles, or they should consult with the Merkle team for consideration.
MeowRPC [58] The transactions are sent to the top builders which MeowRPC’s team trusts.
MEV Blocker [60] Builders must maintain at least 1% market share each week of the previous month.
MEV-Share [37] Builders must demonstrate a competitive market share for consideration.
Transaction Boost [8] Consideration for partnership is based on the Blocknative team’s evaluation of the builder’s reputation.
@@
Table 5: Collections of the private order flow providers and their requirements.
F.5 Minimal Cost for 1% Market Share
We compute the minimal subsidy needed to establish a 1% market share for one month (which is the requirement of MEV Blocker [60]) by replaying historical auctions with an added new builder who can only access the public order flow. We assume that all builders have access to public order flows but the new builder cannot access any of the private order flows. Furthermore, this new builder can win the auction if their bid value is the same as other builders. Other builders will not submit bids beyond their true value, defined as the total revenue of their blocks. Or they will not increase subsidies if they already subsidize. For simplicity of computation, we assume that one month has 30 days.
@@ F.5 Minimal Cost for 1% Market Share
The result is plotted in Fig. 16. Initially, this cost was nearly zero before May 2023, but it has steadily increased, surpassing 7.4 ETH by July 2024.It is important to recognize that the costs derived from this method only represent the lower bound. Predicting the minimum cost of winning a MEV-Boost auction is challenging, so builders need to pay more to ensure victory. Furthermore, builders may subsidize more blocks than required, as it is also difficult to precisely achieve a 1% market share. In reality, the subsidies new builders need to provide are substantially higher. For example, tbuilder [78], a new builder in the builder market, subsidized 6.8 ETH in total for 645 blocks in the week before March 21, 2024 [36].
Figure 16: Subsidy required to establish a 1% market share, as required by MEV Blocker. This quantifies the cost of accessing private order flows.
Appendix G Pseudocode of TEE Builder
The pseudocode for the builder running in TEEs with a refund mechanism is shown in Fig. 17.
[…]
$orders\leftarrow[]$ // Pending transactions.
$sources\leftarrow\{\}$ // Record the order flow providers.
// Receive order flows from providers.
[…]
$orders\leftarrow orders+[order]$
$sources[order]\leftarrow provider$
// Build best block and refund providers.
On building:
$v,B\leftarrow\textsc{build}(orders)$
$payments\leftarrow[]$
// For each provider, refund its marginal contribution,
 For every $order$ in $B$:
  $provider\leftarrow sources[order]$
  $v^{\prime},\_\leftarrow\textsc{build}(orders-[order])$
  $payment\leftarrow\textsc{transfer}(provider,v-v^{\prime})$
  $payments\leftarrow payments+[payment]$
// Remove all transactions included on chain.
 For every $order$ in $B$:
  $orders\leftarrow orders-[order]$
  Delete $sources[order]$
$B\leftarrow B+payments$
[…]
Figure 17: The pseudocode of TEE builder.